OpenStudy (anonymous):

How do I solve this using identity: cos^2x/(1+2sinx-3sin^2x) = (1+sinx)/(1+3sinx)

OpenStudy (anonymous):

let sin (x) be b

OpenStudy (anonymous):

i have to use trig identities tho^

OpenStudy (anonymous):

(1+2sinx-3sin^2x) 1+2b-3 b^2 = (1+3b)(1-b)

OpenStudy (anonymous):

cos^2x use identiry sin^2(x)+cos^2(x)=1 cos^2=1-sin^2(x)---->1-b^2 1-b^2=(1+b)(1-b)

OpenStudy (anonymous):

(1+b)(1-b) -------------- (1+3b)(1-b) (1-b) cancels leaving (1+b) ------------------ (1+3b) remember b=sin(x) 1+

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem